A 35.4 g bullet strikes a 18.0 kg stationary wooden block and embeds itself in the block. The block and the bullet fly off together at 13.3 m/s.



What was the initial velocity of the bullet?
1
Expert's answer
2020-04-15T10:40:29-0400

mass of the bullet "m=35.4" g

mass of the wooden block "M=18" kg

velocity of block and bullet afterwards "v=13.3" m/s

We need to find initial velocity of bullet "v_0"

Let's write conservation law of momentum:

"mv_0=(m+M)v"

Then "v_0=v{\frac {m+M} {m}}=6776" m/s